OpenStudy (anonymous):

simplify (x-2)(9x-8)-(3x-4)^2 @babalooo

OpenStudy (babalooo):

my advise right at this moment is to expand everything

OpenStudy (babalooo):

Okay, here some steps you can take 1) expand (x-2)(9x-8)=solution a 2) expand (3x-4)(3x-4)= solution b 3) solution a- solution b (remember to multiply the neg (-) sign into the signs of b) 4) End up with a final answer.

OpenStudy (babalooo):

Did you expanded it yet?

OpenStudy (anonymous):

Do you mean factor

OpenStudy (gunsodo):

He means that you have to multiply the (x-2) and (9x-8) and multiply (3x-4) by itself, then finally, you subtract the equation by the same degree ones

OpenStudy (babalooo):

No. expand. the opposite of factor then Collect like terms

OpenStudy (babalooo):

Just follow the multiplication equations I've described in step 1 and 2

OpenStudy (gunsodo):

You'll get (x-2)(9x-8)-(3x-4)^2 =(x-2)(9x-8)-(3x-4)(3x-4) =(9x^2-8x-18x+16)-(9x^2-12x-12x+16) =(9x^2-26x+16)-(9x^2-24x+16) =9x^2-26x+16-9x^2+24x-16 =-2x -2x is the final answer...

OpenStudy (gunsodo):

Am I right ?? @babalooo

OpenStudy (babalooo):

@gunsodo yeah, I got that too.. but you're not supposed to do the work for them!!! :D

OpenStudy (babalooo):

lol well i'm gonna go sleep... it's almost 3am. PEACE :'D

OpenStudy (gunsodo):

Nah, I just helped them...

OpenStudy (gunsodo):

Good night, in my country it's just 5 pm. LOL